I havent taken math since college and am helping my Gfriend with her Algerbra........I just dont know how to set up the equation for this word problem. Could you show me how to set it up and solve it?? The question is as follows:
Mr.A owns two appliance repair shops. This year the telephone bill for his Salem store was $610 less than twice the telephone bill for his Saugus shop. The total telephone bill for both stores was $2,054. What was the cost of each shops telephone service for the year?

Answer by Paul(988)

Last year's wage = x
This year's wage = 2x-610
TOgether:

x+2x-610=2054
3x-610+610=2054+610
3x=2664
x=888

2(888)-610
=1166

Hence, this year the wage was $1166.
Paul.